# Large AI Model for Delay-Doppler Domain Channel Prediction in 6G OTFS-Based Vehicular Networks

### v1

@@ III System Model
III-A Large-Scale Fading
@@
$R_{\mathrm{K}}^{\mu}=R_{\mathrm{K}}^{0}+R_{\mathrm{K}}^{\mathrm{f}}\log_{10}\left(f_{\mathrm{G}}\right)+R_{\mathrm{K}}^{\mathrm{d}}\log_{10}\left(d_{2\mathrm{D}}\right)+$ (2)
$R_{\mathrm{K}}^{\mathrm{h}}\log_{10}\left(h_{B}\right)+R_{\mathrm{K}}^{\alpha}\alpha_{\mathrm{R}},$
@@
$L_{\mathrm{P}}=L_{\mathrm{P}}^{0}+L_{\mathrm{P}}^{\mathrm{f}}\log_{10}\left(f_{\mathrm{GHz}}\right)+L_{\mathrm{P}}^{\mathrm{d}}\log_{10}\left(d_{2\mathrm{D}}\right)+$ (6)
$L_{\mathrm{P}}^{\mathrm{h}}\log_{10}\left(h_{B}\right)+L_{\mathrm{P}}^{\alpha}\alpha_{\mathrm{R}},$
@@
III-B Small-Scale Fading
@@
III-C Quasi Deterministic Channel Parameters
@@
IV OTFS and Delay-Doppler Domain Channel
@@ IV-A OTFS Modulation
$\mathbf{Y_{DD}}[l,k]=\frac{1}{\sqrt{MN}}\sum_{n=0}^{N-1}\sum_{m=0}^{M-1}\mathbf{Y_{TF}}[m,n]e^{-j2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}.$ (13)
IV-B Delay-Doppler Domain Channel
@@
$r(t)$ $=\sum_{i=1}^{N_{\mathrm{P}}}{h_{i}}(\tau_{i},t)s\left(t-\tau_{i}\right)$ (16)
$=\sum_{i=1}^{N_{\mathrm{P}}}{g\left(i,t_{0}\right)\cdot e^{j2\pi\nu(i)\left(t-\tau_{i}\right)}}s\left(t-\tau_{i}\right),$
@@
IV-C Channel Prediction Problem Formulation
@@
V Large AI Model for Channel Prediction
@@ V-A Transformer Backbone
V-B Time Series Large AI Model
@@ VI-A Experimental Settings
For the time series large AI model named Timer, the model dimension is configured to 512 and the feed forward dimension is 2048, while it employs 8 heads within the multi-head attention mechanism [19]. During the fine-tuning process, the learning rate is $1\times 10^{-5}$ and the loss function is MAE. Another large AI model named Time-MoE (mixture of experts) is also used for zero-shot prediction [47]. Two classical models including LSTM and GRU are used to predict the channel after training. The input time series length is 20 and the output time series length is 10, indicating that the model predicts future channel conditions based on 10 milliseconds of historical data to forecast the subsequent 5 milliseconds.
VI-B Prediction Accuracy of Delay-Doppler Domain Parameter
TABLE I: Comparison of prediction errors for different velocities in LOS channel scenario.
@@ VI-B Prediction Accuracy of Delay-Doppler Domain Parameter
Timer (zero-shot) 60 2.53E-11 0.0434 0.0267 0.0013
120 1.02E-10 0.1636 0.0493 0.0054
Timer (fine-tuning) 60 7.02E-12 0.0187 0.0101 0.0006
120 3.74E-11 0.0840 0.0256 0.0027
TABLE II: Comparison of prediction errors for different velocities in NLOS channel scenario.
@@
Timer (zero-shot) 60 8.81E-11 0.2499 0.1270 0.0076
120 2.04E-10 0.5884 0.1484 0.0194
Timer (fine-tuning) 60 2.84E-11 0.1277 0.0587 0.0039
120 7.70E-11 0.3299 0.0848 0.0107
Table I and Table II present a comprehensive analysis of different models’ effectiveness in predicting channel conditions in vehicular networks under both LOS and NLOS scenarios. The results highlight the significant advantage of the fine-tuned Timer model, which consistently outperforms classical models and zero-shot large AI models across all metrics with vehicle speeds of 60 km/h and 120 km/h. In LOS conditions, the fine-tuned Timer demonstrates its ability to capture channel dynamics effectively, even as vehicle speed increases, where traditional models like LSTM and GRU struggle to adapt. In NLOS conditions, where channel characteristics become even more complex due to multi-path propagation, the fine-tuned Timer maintains its advantage, highlighting its robustness in handling the severe Doppler shifts and multi-path effects characteristic of NLOS environments. The results also underscore the value of the zero-shot capability inherent in large AI models like Timer and Time-MoE, which enables reasonably accurate channel predictions without specific training, thus facilitating low-cost deployment in new scenarios. Moreover, fine-tuning further refines Timer’s predictive accuracy and enables it to adapt to the unique characteristics of vehicular channels, emphasizing the additional accuracy gained through fine-tuning for task-specific data. In comparison, classical models like LSTM and GRU exhibit considerably higher prediction errors across both LOS and NLOS scenarios, especially as vehicle speed increases, indicating their limitations in capturing the rapid and complex dynamics of vehicular channels.

### v2

# LAIM for Channel Prediction

@@ 3 System model
3.1 Large-scale fading
@@
$R_{\mathrm{K}}^{\mu}=R_{\mathrm{K}}^{0}+R_{\mathrm{K}}^{\mathrm{f}}\log_{10}\left(f_{\mathrm{G}}\right)+R_{\mathrm{K}}^{\mathrm{d}}\log_{10}\left(d_{2\mathrm{D}}\right)+R_{\mathrm{K}}^{\mathrm{h}}\log_{10}\left(h_{B}\right)+R_{\mathrm{K}}^{\alpha}\alpha_{\mathrm{R}},$ (2)
@@
$L_{\mathrm{P}}=L_{\mathrm{P}}^{0}+L_{\mathrm{P}}^{\mathrm{f}}\log_{10}\left(f_{\mathrm{GHz}}\right)+L_{\mathrm{P}}^{\mathrm{d}}\log_{10}\left(d_{2\mathrm{D}}\right)+L_{\mathrm{P}}^{\mathrm{h}}\log_{10}\left(h_{B}\right)+L_{\mathrm{P}}^{\alpha}\alpha_{\mathrm{R}},$ (6)
@@
3.2 Small-scale fading
@@
3.3 Quasi deterministic channel parameters
@@
4 OTFS and delay-Doppler domain channel
@@ 4.1 OTFS modulation
$\mathbf{Y_{DD}}[l,k]=\frac{1}{\sqrt{MN}}\sum_{n=1}^{N}\sum_{m=1}^{M}\mathbf{Y_{TF}}[m,n]e^{-j2\pi\left(\frac{nk}{N}-\frac{ml}{M}\right)}.$ (13)
4.2 Delay-Doppler domain channel
@@
$r(t)=\sum_{i=1}^{N_{\mathrm{P}}}{h_{i}}(\tau_{i},t)s\left(t-\tau_{i}\right)=\sum_{i=1}^{N_{\mathrm{P}}}{g\left(i,t_{0}\right)\cdot e^{j2\pi\nu(i)\left(t-\tau_{i}\right)}}s\left(t-\tau_{i}\right),$ (16)
@@
4.3 Channel prediction problem formulation
@@
5 Large AI model for channel prediction
@@ 5.1 Transformer backbone
5.2 Time series large AI model
@@ 6.1 Experimental settings
For the time series large AI model named Timer, the model dimension is configured to 1024 and the feed forward dimension is 2048, while it employs 8 heads within the multi-head attention mechanism [19]. During the fine-tuning process, the learning rate is $1\times 10^{-5}$ and the loss function is MAE. Another large AI model named Time-MoE (mixture of experts) is also used for zero-shot prediction [47]. Two classical models including LSTM and GRU are used to predict the channel after training. The input time series length is 20 and the output time series length is 10, indicating that the model predicts future channel conditions based on 10 milliseconds of historical data to forecast the subsequent 5 milliseconds.
6.2 Prediction accuracy of delay-Doppler domain parameter
Table 1: Comparison of prediction errors for different velocities in LOS channel scenario.
@@ 6.2 Prediction accuracy of delay-Doppler domain parameter
Timer (zero-shot) 60 1.64E-11 0.0342 0.0202 0.0010
120 5.67E-11 0.2041 0.0986 0.0062
Timer (fine-tuning) 60 6.06E-12 0.0176 0.0081 0.0005
120 2.30E-11 0.1215 0.0519 0.0036
Table 2: Comparison of prediction errors for different velocities in NLOS channel scenario.
@@
Timer (zero-shot) 60 7.26E-11 0.1353 0.0393 0.0043
120 1.43E-10 0.5045 0.1218 0.0160
Timer (fine-tuning) 60 3.14E-11 0.0809 0.0226 0.0026
120 6.33E-11 0.3212 0.0791 0.0104
Table 1 and Table 2 present a comprehensive analysis of different models’ effectiveness in predicting channel conditions in vehicular networks under both LOS and NLOS scenarios. The results highlight the significant advantage of the fine-tuned Timer model, which consistently outperforms classical models and zero-shot large AI models in most metrics with vehicle speeds of 60 km/h and 120 km/h. In LOS conditions, the fine-tuned Timer demonstrates its ability to capture channel dynamics effectively, even as vehicle speed increases, where traditional models like LSTM and GRU struggle to adapt. In NLOS conditions, where channel characteristics become even more complex due to multi-path propagation, the fine-tuned Timer maintains its advantage, highlighting its robustness in handling the severe Doppler shifts and multi-path effects characteristic of NLOS environments. The results also underscore the value of the zero-shot capability inherent in large AI models like Timer and Time-MoE, which enables reasonably accurate channel predictions without specific training, thus facilitating low-cost deployment in new scenarios. Moreover, fine-tuning further refines Timer’s predictive accuracy and enables it to adapt to the unique characteristics of vehicular channels, emphasizing the additional accuracy gained through fine-tuning for task-specific data. In comparison, classical models like LSTM and GRU exhibit considerably higher prediction errors across both LOS and NLOS scenarios, especially as vehicle speed increases, indicating their limitations in capturing the rapid and complex dynamics of vehicular channels.
@@ 7 Conclusion
[…]
This work is supported in part by the National Natural Science Foundation of China (Grant Nos. 623B2052 and 62271244).